OpenStudy (anonymous):

find solutions for x: \[\sqrt{x^-11=5}\]

OpenStudy (slaaibak):

you sure that's the question?

OpenStudy (anonymous):

\[\sqrt{x-11}=5\] ?

OpenStudy (anonymous):

yes that is it

OpenStudy (slaaibak):

36

OpenStudy (anonymous):

square both sides and add 11. x-11=25 x=36

OpenStudy (anonymous):

the x is squared

OpenStudy (anonymous):

the x is squared so its x2